Helen invested $10,000 in an account that pays 10% simple interest. How much additional money must be invested in an account that pays 13% simple interest so that the average return on the two investments amounts to 11%?
Answer by HyperBrain(694)

let x=amount invested at 13%
then
x+10000=total money invested
(0.10)10000+0.13x=0.11(x+10000)
1000+0.13x=0.11x+1100
0.13x=0.11x+100
0.02x=100
2x=10000
x=$5000
To get an average 11% return, Helen must invest $5000 at 13%
Power up,
HyperBrain!
visit my bloghere